# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 3.3 Approximate solution
A partition of Kerr de Sitter is not a solution to the Einstein equations. However, all that is needed for our purposes is that it is an approximate solution. It turns out that the following is sufficient.
@@
The partition metric (3.12) satisfies:
$\left\{\begin{array}[]{ccc}\big{|}\partial_{x}^{\alpha}\big{(}\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda\big{)}\big{|}&\leq\quad C_{\alpha}e^{-2Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\\ \big{|}\partial_{x}^{\alpha}\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}\big{|}&\leq\quad C_{\alpha}e^{-4Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\\ \big{|}\partial_{x}^{\alpha}\big{(}\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}\big{)}\big{|}&\leq\quad C_{\alpha}e^{-2Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\end{array}\right.,\qquad\text{for $|t|<1$},$ (3.19)
and
$\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda=\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}=\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}=0,\qquad\text{for $|t|\geq 1$}.$ (3.20)
@@
${\bm{\nabla}}_{\partial_{s}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}{\nabla}\widetilde{\Phi}$
$\bm{\widetilde{\Gamma}}_{ss}^{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\qquad\bm{\widetilde{\Gamma}}_{ss}^{i}=\widetilde{\Phi}\widetilde{g}^{ij}\partial_{j}\widetilde{\Phi}=\mathcal{O}(e^{-4Hs})$
Moreover, we know ${\bm{\nabla}}_{\partial_{i}}N=k_{i}{}^{j}\partial_{j}$
${\bm{\nabla}}_{\partial_{i}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\partial_{j}$
$\bm{\widetilde{\Gamma}}_{is}^{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\qquad\bm{\widetilde{\Gamma}}_{is}^{j}=\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})$
@@
$\bm{\widetilde{\Gamma}}_{ij}^{s}=-\widetilde{\Phi}^{-1}\widetilde{{\bm{g}}}({\bm{\nabla}}_{\partial_{i}}\partial_{j},\widetilde{\Phi}^{-1}\partial_{s})=\widetilde{\Phi}^{-1}\widetilde{k}_{ij}=HG_{ij}(x)e^{2Hs}+\mathcal{O}(1)$
@@
$\widetilde{\Gamma}_{ij}^{k}=\frac{1}{2}G^{kl}\bigl{(}\partial_{i}G_{jl}+\partial_{j}G_{il}-\partial_{l}G_{ij}\bigr{)}+\mathcal{O}(e^{-2Hs})$
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\alpha}=$ $\mathcal{O}(e^{-2Hs})\,,\qquad\partial_{s}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,,$
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{js}$ $=\mathcal{O}(e^{-2Hs})\,,\qquad\partial_{j}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,,$
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{ji}$ $=2H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)\,,\qquad\partial_{j}{\bm{\widetilde{\Gamma}}}^{\alpha}_{\alpha i}=\mathcal{O}(1)\,,$
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{ij}=3H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)=\Lambda\widetilde{g}_{ij}+\mathcal{O}(1)$ (3.23)
@@
$\bm{\widetilde{\Gamma}}_{ss}^{s}=$ $\,\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}={{}^{\mathcal{K}}\bm{\Gamma}}_{ss}^{s}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ss}^{i}=$ $\,\widetilde{\Phi}\widetilde{g}^{ij}\partial_{j}\widetilde{\Phi}={}^{\mathcal{K}}\bm{\Gamma}_{ss}^{i}+e^{-2Hs}\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{is}^{s}=$ $\,\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}={}^{\mathcal{K}}\bm{\Gamma}_{is}^{s}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{is}^{j}=$ $\,\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=\Phi_{\mathcal{K}_{a,m}}(k_{\mathcal{K}_{a,m}})_{i}{}^{j}+\mathcal{O}(\widetilde{\varepsilon})={}^{\mathcal{K}}\bm{\Gamma}_{is}^{j}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ij}^{s}=$ $\,\widetilde{\Phi}^{-1}\widetilde{k}_{ij}={}^{\mathcal{K}}\bm{\Gamma}_{ij}^{s}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ij}^{k}=$ $\,{}^{\mathcal{K}}\bm{\Gamma}_{ij}^{k}+\mathcal{O}(\widetilde{\varepsilon})$
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{ss}=\partial_{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\alpha}-\partial_{s}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}+\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\gamma}-\bm{\widetilde{\Gamma}}_{s\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}\\ ={}^{\mathcal{K}}{\bm{\mathrm{Ric}}}_{ss}+\mathcal{O}(\widetilde{\varepsilon})=\Lambda({\bm{g}}_{\mathcal{K}_{a,m}})_{ss}+\mathcal{O}(\widetilde{\varepsilon})=\Lambda\widetilde{{\bm{g}}}_{ss}+\mathcal{O}(\widetilde{\varepsilon}),$
because $\widetilde{{\bm{g}}}_{ss}=-\widetilde{\Phi}^{2}=-\Phi_{\mathcal{K}_{a,m}}^{2}+\mathcal{O}(\widetilde{\varepsilon})$. Hence, $\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\Lambda=\mathcal{O}(\widetilde{\varepsilon})$.
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{sj}=\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{js}-\partial_{j}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}+\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{js}^{\gamma}-\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}={}^{\mathcal{K}}{\bm{\mathrm{Ric}}}_{sj}+\mathcal{O}(\widetilde{\varepsilon})=\mathcal{O}(\widetilde{\varepsilon})$
where we have used that $\bm{\widetilde{\Gamma}}_{ji}^{s}\bm{\widetilde{\Gamma}}_{ss}^{i}=\mathcal{O}(\widetilde{\varepsilon})$.
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{ji}$ $=\partial_{\alpha}{}^{\mathcal{K}}\bm{\Gamma}^{\alpha}_{ji}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,,\qquad\partial_{j}{\bm{\widetilde{\Gamma}}}^{\alpha}_{\alpha i}=\partial_{j}{}^{\mathcal{K}}\bm{\Gamma}^{\alpha}_{\alpha i}+\mathcal{O}(\widetilde{\varepsilon})\,,$
$\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{ji}^{\gamma}$ $={}^{\mathcal{K}}\bm{\Gamma}_{\alpha\gamma}^{\alpha}{}^{\mathcal{K}}\bm{\Gamma}_{ji}^{\gamma}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,,\qquad\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{\alpha i}^{\gamma}={}^{\mathcal{K}}\bm{\Gamma}_{j\gamma}^{\alpha}{}^{\mathcal{K}}\bm{\Gamma}_{\alpha i}^{\gamma}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,.$
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{ij}={\bm{\mathrm{Ric}}}[{\bm{g}}_{\mathcal{K}_{a,m}}]+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})=\Lambda\widetilde{{\bm{g}}}_{ij}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})$
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}=\mathcal{O}(\widetilde{\varepsilon})\,.$
@@ 4.1 Weighted norms and energy
For the spacetimes we consider, $\Sigma_{s}$ is diffeomorphic to ×$\mathbb{R}\times\mathbb{S}^{2}$. While $g=g_{s}$ is a metric on $\Sigma_{s}$, we endow (×$(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ with the standard metric $\mathring{g}$ on the cylinder:
$\mathring{g}=\mathrm{d}t^{2}+\mathring{\gamma},\qquad\mathring{\gamma}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}.$ (4.1)
The coordinate charts covering $\Sigma_{s}$ are denoted by $(t,\theta,\phi)$; recall that by construction these are coordinates on each $\Sigma_{s}$, $s\geq s_{0}$, and can also be identified with coordinates on a chart for ×$\mathbb{R}\times\mathbb{S}^{2}$. Given a $\Sigma_{s}$-tangent $(n,m)$ tensor $\mathcal{T}$, we define
@@
We define ̊$\mathring{\nabla}^{(\ell)}\mathcal{T}$ and $|\mathcal{T}|_{\mathring{g}}^{2}$ similarly, using the covariant derivative $\mathring{\nabla}$ of $\mathring{g}$, instead of $\nabla$, and contracting indices with $\mathring{g}$ instead of $g$.
@@
where $\mathrm{vol}_{g}=\sqrt{|g|}\>\mathrm{d}t\wedge\mathrm{d}\theta\wedge\mathrm{d}\phi$ is the volume form of $(\Sigma_{s},g)$ and
@@
The exponential rates $\alpha_{1},\alpha_{2}$ are related to the exponential decay of the perturbed solution towards the endpoints $t=\pm\infty$. They are nonnegative, and may not be equal; they can also be set to 00, when no exponential decay along the cosmological horizons is imposed initially. While the norm $\|\cdot\|_{H^{M}}$ does depend on $\alpha_{1},\alpha_{2}$, we typically supress this in the notation. To simplify notation, we often drop $(\Sigma_{s},g)$ from subscript to the norms (4.3), (4.4), (4.5), and (4.7).
@@
$\begin{split}\mathcal{E}_{N}(s)=&\,\|\widehat{g}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{H^{N}(\Sigma_{s},g)}\\ &+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{H^{N}(\Sigma_{s},g)}\big{)}\end{split}$ (4.11)
@@ 4.2 Bootstrap assumptions and basic consequences
Our bootstrap assumptions are that there exists a bootstrap time $s_{b}\in(s_{0},+\infty)$ such that the following inequalities hold:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{2Hs},\qquad\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{-2Hs},\qquad\mathcal{E}_{N}(s)\leq\varepsilon^{2},$ (4.12)
for all $s\in[s_{0},s_{b})$, and some $N\geq 4$. Notice that such a bootstrap time exists from classical Cauchy stability.
@@ 4.3 Preliminary estimates
$C^{-1}|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|\leq e^{(m-n)Hs}|\mathcal{T}|_{g}\leq C|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|$ (4.13)
for all $i_{1},\ldots,i_{m},j_{1},\ldots,j_{n}$ and $(s,x)\in(s_{0},s_{b})\times\Sigma_{s}$.
@@
The classical Sobolev embedding in ×$\mathbb{R}\times\mathbb{S}^{2}$ implies that
@@
since |̊$|\mathring{\nabla}^{(\ell)}f(t)|\leq Cf(t)$. Recall (4.10) and use Lemma 4.7 to deduce that
@@
$\nabla(\nabla(\mathcal{T}))=$ $\,\nabla\nabla\mathcal{T}+\Gamma\star\nabla\mathcal{T}+\Gamma\star\Gamma\star\mathcal{T}+\nabla\Gamma\star\mathcal{T}$

### v2

@@ 3.3 Approximate solution
A partition of Kerr de Sitter is not a solution to the Einstein equations (in the transition region $|t|<1$). However, all that is needed for our purposes is that it is an approximate solution. It turns out that the following is sufficient.
@@
The partition metric (3.13) satisfies, on one hand:
$\left\{\begin{array}[]{ccc}\big|\partial_{x}^{\alpha}\big(\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda\big)\big|&\leq\qquad C_{\alpha}e^{-2Hs}\\ \big|\partial_{x}^{\alpha}\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}\big|&\leq\qquad C_{\alpha}e^{-4Hs}\\ \big|\partial_{x}^{\alpha}\big(\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}\big)\big|&\leq\qquad C_{\alpha}e^{-2Hs}\end{array}\right.,\hskip 18.49988pt\text{for $|t|<1$},$ (3.20)
and on the other:
$\big|\partial_{x}^{\alpha}\big(\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda\big)\big|+\big|\partial_{x}^{\alpha}\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}\big|+\big|\partial_{x}^{\alpha}\big(\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}\big)\big|\leq C_{\alpha}\widetilde{\varepsilon}\,,\qquad\text{for $|t|<1$}.$ (3.21)
Moreover,
$\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda=\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}=\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}=0,\hskip 18.49988pt\text{for $|t|\geq 1$}.$ (3.22)
@@
${\bm{\nabla}}_{\partial_{s}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}{\nabla}\widetilde{\Phi}\,$,
$\bm{\widetilde{\Gamma}}_{ss}^{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\hskip 18.49988pt\bm{\widetilde{\Gamma}}_{ss}^{i}=\widetilde{\Phi}\widetilde{g}^{ij}\partial_{j}\widetilde{\Phi}=\mathcal{O}(e^{-4Hs})\,.$
Moreover, we know ${\bm{\nabla}}_{\partial_{i}}N=k_{i}{}^{j}\partial_{j}$,
${\bm{\nabla}}_{\partial_{i}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\partial_{j}\,$,
$\bm{\widetilde{\Gamma}}_{is}^{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\hskip 18.49988pt\bm{\widetilde{\Gamma}}_{is}^{j}=\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})\,.$
@@
$\bm{\widetilde{\Gamma}}_{ij}^{s}=-\widetilde{\Phi}^{-1}\widetilde{{\bm{g}}}({\bm{\nabla}}_{\partial_{i}}\partial_{j},\widetilde{\Phi}^{-1}\partial_{s})=\widetilde{\Phi}^{-1}\widetilde{k}_{ij}=HG_{ij}(x)e^{2Hs}+\mathcal{O}(1)\,$,
@@
$\widetilde{\Gamma}_{ij}^{k}=\frac{1}{2}G^{kl}\bigl(\partial_{i}G_{jl}+\partial_{j}G_{il}-\partial_{l}G_{ij}\bigr)+\mathcal{O}(e^{-2Hs})\,.$
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\alpha}=$ $\mathcal{O}(e^{-2Hs})\,,\hskip 18.49988pt\partial_{s}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,$,
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{js}$ $=\mathcal{O}(e^{-2Hs})\,,\hskip 18.49988pt\partial_{j}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,$,
@@
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{ji}$ $=2H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)\,,\hskip 18.49988pt\partial_{j}{\bm{\widetilde{\Gamma}}}^{\alpha}_{\alpha i}=\mathcal{O}(1)\,$,
@@
$\widetilde{{\bm{\mathrm{Ric}}}}_{ij}=3H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)=\Lambda\widetilde{g}_{ij}+\mathcal{O}(1)\,$, (3.25)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 The bootstrap argument
In this section we prove the global existence statement of Theorem 1.2 (I).
@@ 4.1 Weighted norms and energy
For the spacetimes we consider, $\Sigma_{s}$ is diffeomorphic to $\mathbb{R}\times\mathbb{S}^{2}$. While $g=g_{s}$ is a metric on $\Sigma_{s}$, we endow $(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ with the standard metric $\mathring{g}$ on the cylinder:
$\mathring{g}=\mathrm{d}t^{2}+\mathring{\gamma},\hskip 18.49988pt\mathring{\gamma}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}.$ (4.1)
The coordinates on $\Sigma_{s}$ are denoted by $(t,\theta,\phi)$; recall that by construction these are coordinates on each $\Sigma_{s}$, $s\geq s_{0}$, and can also be identified with coordinates on a chart for $\mathbb{R}\times\mathbb{S}^{2}$. Given a $\Sigma_{s}$-tangent $(n,m)$ tensor $\mathcal{T}$, we define
@@
[…]
@@
where $\mathrm{vol}_{g}=\sqrt{|g|}\>\mathrm{d}t\wedge\mathrm{d}\theta\wedge\mathrm{d}\phi$ is the volume form of $(\Sigma_{s},g)$ in local coordinates and
@@
The exponential rates $\alpha_{1},\alpha_{2}$ are related to the exponential decay of the perturbed solution towards the endpoints $t=\pm\infty$. They are nonnegative, and may not be equal; they can also be set to 0, when no exponential decay along the cosmological horizons is imposed initially. While the norm $\|\cdot\|_{H^{M}}$ does depend on $\alpha_{1},\alpha_{2}$, we typically supress this in the notation. To simplify notation, we often drop $(\Sigma_{s},g)$ from subscript to the norms (4.3), (4.4), (4.5), and (4.7).
@@
[…]
@@ 4.2 Bootstrap assumptions and basic consequences
In view of the local well-posedness statement in Appendix A, there exists $\varepsilon>0$ and a maximal time of existence $s_{b}\in(s_{0},+\infty)$ such that the following inequalities hold:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{2Hs},\hskip 18.49988pt\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{-2Hs},\hskip 18.49988pt\mathcal{E}_{N}(s)\leq\varepsilon^{2},$ (4.12)
for all $s\in[s_{0},s_{b})$, and some $N\geq 4$.
These are our bootstrap assumptions, and the statement that such a bootstrap time $s_{b}\in(s_{0},+\infty)$ exists follows from classical Cauchy stability for the locally well-posed system of Appendix A & B. In Section 4.4 we will show that for data sufficiently close to Kerr de Sitter, these bounds are not saturated at $s=s_{b}$, and the solution can hence be continued, using again the local existence statement of Appendix A, to yield the global existence result of Theorem 1.2 (I).
@@ 4.3 Preliminary estimates
$C^{-1}\min_{\begin{subarray}{c}i_{1},\ldots,i_{m},\\ j_{1},\ldots,j_{n}\end{subarray}}|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|\leq e^{(m-n)Hs}|\mathcal{T}|_{g}\leq C\max_{\begin{subarray}{c}i_{1},\ldots,i_{m},\\ j_{1},\ldots,j_{n}\end{subarray}}|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|$ (4.13)
for all $(s,x)\in(s_{0},s_{b})\times\Sigma_{s}$.
@@
The classical Sobolev embedding in $\mathbb{R}\times\mathbb{S}^{2}$ implies that
@@
since $|\mathring{\nabla}^{(\ell)}f(t)|\leq Cf(t)$. Recall (4.10) and use Lemma 4.7 to deduce that
@@
$\nabla(\nabla(\mathcal{T}))=$ $\,\nabla\nabla\mathcal{T}+\Gamma\star\nabla\mathcal{T}+\Gamma\star\Gamma\star\mathcal{T}+\nabla\Gamma\star\mathcal{T}\,.$